For my assignment: Suppose that the length of a rectangle is three cm longer than twice the width and that the perimeter of the rectangle is 90.
1. set up an equation for the perimeter involving only W, the width of the rectangle.

Answer by htmentor(1343)

For my assignment: Suppose that the length of a rectangle is three cm longer than twice the width and that the perimeter of the rectangle is 90.
1. set up an equation for the perimeter involving only W, the width of the rectangle.
================================
Let l = the length, w = the width
Given:
length is 3 cm MORE than TWICE the width
l = 3 + 2w
perimeter is 90.
2l + 2w = 90
Substitute the value for l from the 1st equation into the 2nd:
2(3+2w) + 2w = 90
Simplify:
6 + 4w + 2w = 90
6w = 84
w = 14